(a). The t statistic for a test of

?0:?=20

??:?<20
based on n = 10 observations has the value t = -1.32.

Using the appropriate table in your formula packet, bound the p-value as closely as possible:
_ < p-value < _

(b).

The t statistic for a test of

?0:?=41

??:?>41
based on n = 17 observations has the value t = 1.72.

Using the appropriate table in your formula packet, bound the p-value as closely as possible:
_< p-value <_

(c).

The t statistic for a test of

?0:?=31

??:?≠31
based on n = 6 observations has the value t = -1.1.

Using the appropriate table in your formula packet, bound the p-value as closely as possible:
_ < p-value < _
Answer #1

1). \mathbf{0.10<p-value <0.15}

[ explanation:-

this is a left tailed test.

df = (n-1) = (10-1) = 9

using t distribution table for df=9, t score = -1.32, one tailed test we have:-

0.10 < p value < 0.15 ]

2). \mathbf{0.05<p-value <0.10}

[ explanation:-

this is a right tailed test.

df = (n-1) = (17-1) = 16

using t distribution table for df=16, t score = 1.72, one tailed test we have:-

0.05 < p value < 0.10 ]

3). \mathbf{0.30<p-value <0.40}

[ explanation:-

this is a both tailed test.

df = (n-1) = (6-1) = 5

using t distribution table for df=5, t score = -1.1, both tailed test we have:-

0.30 < p value < 0.40 ]

*** if you have any doubts about the problem please mention it in the comment box... if satisfied please LIKE.